Calculate the molality and van't Hoff factor (i) for the following aqueous solution: 0.775 mass % KCl, freezing point = −0.364°C m = m KCl i =
Answer #1

0.775 mass % means mass of KCl = 0.775 g for 100 g of solution

mass of KCl = 0.775 g

molar mass of KCl = 74.55 g / mole

moles of KCl = 0.775 / 74.55

                     = 0.0104

mass of solvent = 100 - 0.775 = 99.225 g

molality = moles x 1000 / mass of solvent

            = 0.0104 x 1000 / 99.225

           = 0.105 m

molality = 0.105 m

freezing point depression = i x Kf x molality

0.364 = i x 1.86 x 0.105

i = 1.86

van't Hoff factor (i)   = 1.86